# Maximum Degree-Based Quasi-Clique Search via an Iterative Framework

### v1

@@ 3. Our Basic Iterative Framework
  • get-k$(x):=\lfloor(1-\gamma)\cdot(x-1)\rfloor+1$, which takes a number $x$ of vertices as input and returns a appropriate value of $k$;
  • solve-plex$(y):=\text{the size of the largest }y\text{-plex in }G$, which takes a value $y$ as input.
@@ 7      $k_{i+1}\leftarrow\texttt{get-k}(s_{i})$; $i\leftarrow i+1$;
[…]
@@
[…]
@@ 1 $lb,ub\leftarrow\texttt{Get-Bounds}(G,\gamma)$ ;
3while $\min_{v\in V}d_{G}(v)\leq\lfloor(lb-1)\cdot\gamma\rfloor$ do
@@ 4.2. Our Preprocessing Technique
Based on the $k$-core, a related concept is the core number of a vertex $u$, denoted as ${core}(u)$, which is the largest $k$ such that $u$ is part of a $k$-core in the graph. In other words, $u$ cannot belong to any subgraph where all vertices have a degree of at least $core(u)+1$. Let $max\_core$ be the maximum core number in the graph, i.e., $max\_core=\max_{u\in V}core(u)$. Given a graph $G=(V,E)$, the core numbers for all vertices can be computed in $O(|V|+|E|)$ time using the famous peeling algorithm (Matula and Beck, 1983; Batagelj and Zaveršnik, 2003), which iteratively removes vertices with the smallest degree from the current graph. We remark that during the peeling algorithm used to compute the core numbers, we can obtain the lower bound $lb$ of $s^{*}$ as a by-product. In particular, we check whether the current graph qualifies as a $\gamma$-quasi clique and record the size of the largest qualifying $\gamma$-quasi clique. Moreover, as the current graph keeps shrinking during the peeling algorithm, the lower bound $lb$ clearly corresponds to the size of the first current graph that meets the $\gamma$-quasi clique requirements.
@@
The proof of Lemma 4.2 directly follows from the fact that the minimum degree of a $\gamma$-QC is at most $max\_core$ and the definition of $\gamma$-QC.
Output: The lower and upper bounds $lb$ and $ub$ of $s^{*}$
@@ 11      Remove $u$ from $G$;
13return $lb$, $ub$;
Algorithm 5 Get-Bounds$(G,\gamma)$
@@
The preprocessing stage. After obtaining both lower and upper bounds via Get-Bounds, we show how to utilize these bounds in our preprocessing. Recall that, with $lb$ and $ub$, we can (1) reduce the graph by removing unpromising vertices from $G$ that cannot appear in $g^{*}$ using $lb$, and (2) set a smaller initial value of $k$ in the function Plex-Search using $ub$. The details are as follows. For (1), we can iteratively remove vertices from $G$ with degrees at most $\lfloor(lb-1)\cdot\gamma\rfloor$ and their incident edges, since it is clear that no solution of size greater than $lb$ can still include these vertices. For (2), we can easily derive from Lemma A.1 (Section 4 where the proofs are in Appendix A.2) that, instead of initializing $s_{0}$ as $|V|$, $s_{0}$ can be set to an integer larger than $s^{*}$ while still ensuring our iterative framework produces the correct solution. This allows us to use the obtained upper bound $ub$ to update $s_{0}$. Additionally, this $ub$ is used to generate the initial value of $k$ for Plex-Search through get-k, while maintaining the correctness of the iterative framework.
@@ 5.1. Comparison with Baselines
Performance on representative instances. The runtime performance comparison between IterQC and the two baseline algorithms with $\gamma=0.75$ across 30 representative instances is shown in Table 2. As illustrated in the table, IterQC consistently demonstrates superior efficiency, outperforming both baselines FastQC and DDA across nearly all instances. In particular, IterQC can solve all the graph instances, while both baseline algorithms FastQC and DDA exhibit a high occurrence of timeouts, failing to yield solutions within the 3-hour limit. Specifically, FastQC and DDA fail to solve 23 and 16 instances, respectively. Furthermore, IterQC successfully solves 14 out of the 30 representative instances where both baseline algorithms exceed the 3-hour time limit. For example, on G3, IterQC uses only 0.37 second, while both baselines cannot complete in 3 hours, achieving at least a 29,000$\times$ speed-up. These results further suggest the efficiency superiority of IterQC over both baseline algorithms. The superior performance of IterQC, particularly compared to FastQC, may be due to the hereditary property of the $k$-plex, which allows for more efficient pruning during the branch-and-bound search. However, in rare cases, the computational overhead of IterQC exceeds that of DDA, such as in G19, G23, and G26. This is because DDA adopts an iterative approach based on the IP solver CPLEX, which differs fundamentally from the branch-and-bound-based approaches used by IterQC and FastQC. As a general-purpose solver, CPLEX is not specifically optimized for the $\gamma$-quasi-clique problem and is difficult to tailor for it. Consequently, while DDA may occasionally perform better in specific instances, IterQC generally outperforms DDA in most cases.
@@ 5.2. Ablation Studies
Effectiveness of the preprocessing technique. From Table 3, we observe that IterQC consistently outperforms IterQC-PP, achieving a speedup factor of at least 5 in 17 instances and at least 10 in 6 instances, with a remarkable speedup factor of 20.32 on G29. We also summarize additional information for our preprocessing technique in Table 4, which details the percentages of vertices and edges pruned during preprocessing, as well as the lower and upper bounds ($lb$ and $ub$) for the optimum solution $s^{*}$. From Table 4, we observe that in 3 instances (G17, G20, and G29), the preprocessing technique prunes all vertices and edges, effectively obtaining the solution directly, while in 15 instances, it removes at least 90% of the vertices. Moreover, across the 30 instances, the preprocessing technique enables the iteration process to start from a smaller initial value (closer to the optimum solution $s^{*}$), as indicated by the upper bound $ub$ (in contrast to the trivial upper bound $|V|$ in Table 1).
@@ A.1. Omitted Proofs in Section 3
[…]
@@
Assume, to the contrary, that $s_{i}=s_{i+1}$ for $10\leq i\leq p-$ 1. Then, it follows that $k_{i+1}=\texttt{get-k}(s_{i})=\texttt{get-k}(s_{i+1})$, which satisfies the termination condition in Line 4. This implies that the algorithm terminates at the $(i+1)$-st iteration, thus $i=p$, leading to a contradiction to $i\leq p-1$. ∎
@@
[…]
@@ A.2. Correctness Proof of Algorithm 4
Consider the $i$-th iteration of Lines 2-6 in Algorithm 4, where Plex-Search$(G,k,ub\mbox{-}plex)$ is called in Line 3. We let $\overline{s}\leftarrow\texttt{solve-plex}(k)$, i.e., $\overline{s}$ is the size of maximum $k$-plex. If we have $k=\texttt{get-k}(ub\mbox{-}plex)$ and $ub\mbox{-}plex=s_{i-1}\geq s^{*}$, then $s_{i}\geq\overline{s}\geq s^{*}$.
@@
Consider the $i$-th iteration in Lines 2-6 of Algorithm 4, where Plex-Search$(G,k,ub\mbox{-}plex)$ is called in Line 3. Note that $ub\mbox{-}plex=s_{i-1}$. We focus on this Plex-Search and have two cases.
@@ Appendix C Additional Experiments
Ablation studies. Tables 7 to 8 present the results of the ablation studies conducted on 30 representative instances, demonstrating the impact of preprocessing and the pseudo lower bound (pseudo LB) techniques on practical performance. The results clearly indicate significant improvements across all values of $\gamma$. For the preprocessing technique, the optimization effect is clear in nearly every case. For example, at $\gamma=0.65$, instance G29 achieves a speedup of approximately 20.79×20.79$\times$20.79 × compared to the setting without the preprocessing technique. Regarding the pseudo LB technique, substantial benefits are observed in most instances, particularly for smaller values of $\gamma$. At $\gamma=0.65$, 23 out of 30 instances show accelerated running time. Specifically, for G11 at $\gamma=0.65$ and G15 at $\gamma=0.85$, the introduction of the pseudo LB technique transforms previously unsolvable instances within the 3-hour limit into solvable ones. For G15 at $\gamma=0.85$, the running time is reduced to 57.88 seconds, corresponding to a speedup of 186×186$\times$186 ×. These findings highlight the critical contributions of both preprocessing and pseudo LB techniques in improving efficiency of the proposed method.

### v2

@@ 3. Our Basic Iterative Framework
[…]
@@ 7 $k_{i+1}\leftarrow\texttt{get-k}(s_{i})$; $i\leftarrow i+1$;
$s_{p}$ $=\texttt{solve-plex}(k_{p})=\texttt{solve-plex}(\texttt{get-k}(s_{p-1}))$
$\leq\texttt{solve-plex}(\texttt{get-k}(s_{p-2}))=\texttt{solve-plex}(k_{p-1})=s_{p-1}.$
@@
We prove by contradiction. By Lemma 3.6, we assume, to the contrary, that $s_{p+1}>s^{*}$. According to the termination condition in Line 4 of Algorithm 1, we have $k_{p+1}=\texttt{get-k}(s_{p+1})=1+\lfloor(1-\gamma)\cdot(s_{p+1}-1)\rfloor$. Thus, it is clear that the result obtained by $\texttt{solve-plex}(k_{p+1})$ is a $\gamma$-quasi-clique of size $s_{p+1}$, which contradicts $s^{*}$ being the size of the maximum $\gamma$-quasi-clique. ∎
@@ 1 $lb,ub\leftarrow\texttt{Get-Bounds}(G,\gamma)$ ;
3while $\min_{v\in V}d_{G}(v)<\lfloor(lb-1)\cdot\gamma\rfloor$ do
@@ 4.2. Our Preprocessing Technique
Based on the $k$-core, a related concept is the core number of a vertex $u$, denoted as core(u){core}(u), which is the largest $k$ such that $u$ is part of a $k$-core in the graph. In other words, $u$ cannot belong to any subgraph where all vertices have a degree of at least $core(u)+1$. Let max_coremax\_core be the maximum core number in the graph, i.e., $max\_core=\max_{u\in V}core(u)$. Given a graph $G=(V,E)$, the core numbers for all vertices can be computed in $O(|V|+|E|)$ time using the famous peeling algorithm (Matula and Beck, 1983; Batagelj and Zaveršnik, 2003), which iteratively removes vertices with the smallest degree from the current graph. We remark that during the peeling algorithm used to compute the core numbers, we can obtain the lower bound $lb$ of $s^{*}$ as a by-product. In particular, we check whether the current graph qualifies as a $\gamma$-quasi clique and record the size of the largest qualifying $\gamma$-quasi clique. Moreover, as the current graph keeps shrinking during the peeling algorithm, the lower bound $lb$ clearly corresponds to the size of the first current graph that meets the $\gamma$-quasi clique requirements.
@@
The proof of Lemma 4.2 directly follows from the fact that the minimum degree of a $\gamma$-QC is at most max_coremax\_core and the definition of $\gamma$-QC.
Output: The lower and upper bounds lblb and ubub of $s^{*}$
@@ 11  Remove $u$ from $G$;
13return lblb, ubub;
[…]
@@
The preprocessing stage. After obtaining both lower and upper bounds via Get-Bounds, we show how to utilize these bounds in our preprocessing. Recall that, with $lb$ and $ub$, we can (1) reduce the graph by removing unpromising vertices from $G$ that cannot appear in $g^{*}$ using $lb$, and (2) set a smaller initial value of $k$ in the function Plex-Search using $ub$. The details are as follows. For (1), we can iteratively remove vertices from $G$ with degrees less than $\lfloor(lb-1)\cdot\gamma\rfloor$ and their incident edges, since it is clear that no solution of size greater than $lb$ can still include these vertices. For (2), we can easily derive from Lemma A.1 (Section 4 where the proofs are in Appendix A.2) that, instead of initializing $s_{0}$ as $|V|$, $s_{0}$ can be set to an integer larger than $s^{*}$ while still ensuring our iterative framework produces the correct solution. This allows us to use the obtained upper bound $ub$ to update $s_{0}$. Additionally, this $ub$ is used to generate the initial value of $k$ for Plex-Search through get-k, while maintaining the correctness of the iterative framework.
@@ 5.1. Comparison with Baselines
[…]
@@ 5.2. Ablation Studies
Effectiveness of the preprocessing technique. From Table 3, we observe that IterQC consistently outperforms IterQC-PP, achieving a speedup factor of at least 5 in 17 instances and at least 10 in 6 instances, with a remarkable speedup factor of 20.32 on G29. We also summarize additional information for our preprocessing technique in Table 4, which details the percentages of vertices and edges pruned during preprocessing, as well as the lower and upper bounds ($lb$ and $ub$) for the optimum solution $s^{*}$. From Table 4, we observe that in 3 instances (G17, G20, and G29), the preprocessing technique prunes all vertices and edges, effectively obtaining the solution directly, while in 14 instances, it removes at least 90% of the vertices. Moreover, across the 30 instances, the preprocessing technique enables the iteration process to start from a smaller initial value (closer to the optimum solution $s^{*}$), as indicated by the upper bound $ub$ (in contrast to the trivial upper bound $|V|$ in Table 1).
@@ A.1. Omitted Proofs in Section 3
When $G$ is a $\gamma$-quasi-clique, we have $d_{G}(v)\geq\gamma\cdot(|V|-1)$, $\forall v\in V$. Since $d_{G}(v)$ is an integer for each $v\in V$, it follows that $d_{G}(v)\geq\lceil\gamma\cdot(|V|-1)\rceil\geq|V|-(\lfloor(1-\gamma)\cdot(|V|-1)\rfloor+1)=|V|-k_{1}$. Thus, $G$ is a $k_{1}$-plex and $\texttt{solve-k}(k_{1})=|V|$, which implies that $s_{1}=s_{0}$. At this point, we have $k_{1}=\texttt{get-k}(s_{1})$, and the algorithm results in $s_{1}=|V|$, which completes the proof. ∎
@@
Assume, to the contrary, that $s_{i}=s_{i+1}$ for $0\leq i\leq p-1$. Then, it follows that $k_{i+1}=\texttt{get-k}(s_{i})=\texttt{get-k}(s_{i+1})$, which satisfies the termination condition in Line 4. This implies that the algorithm terminates at the $(i+1)$-st iteration, thus $i=p$, leading to a contradiction to $i\leq p-1$. ∎
@@
$s_{i}$ $=\texttt{solve-plex}(k_{i})=\texttt{solve-plex}(\texttt{get-k}(s_{i-1}))$
$=\texttt{solve-plex}\left(1+\left\lfloor(1-\gamma)\cdot(s_{i-1}-1)\right\rfloor\right)$
$\geq\texttt{solve-plex}\left(1+\left\lfloor(1-\gamma^{*})\cdot(s_{i-1}-1)\right\rfloor\right)$
$\geq\texttt{solve-plex}\left(1+\left\lfloor(1-\gamma^{*})\cdot(s^{*}-1)\right\rfloor\right)=s^{*}.$
@@ A.2. Correctness Proof of Algorithm 4
[…]
@@
[…]
@@ Appendix C Additional Experiments
Ablation studies. Tables 7 to 8 present the results of the ablation studies conducted on 30 representative instances, demonstrating the impact of preprocessing and the pseudo lower bound (pseudo LB) techniques on practical performance. The results clearly indicate significant improvements across all values of $\gamma$. For the preprocessing technique, the optimization effect is clear in nearly every case. For example, at $\gamma=0.65$, instance G29 achieves a speedup of approximately $20.79\times$ compared to the setting without the preprocessing technique. Regarding the pseudo LB technique, substantial benefits are observed in most instances, particularly for smaller values of $\gamma$. At $\gamma=0.65$, 23 out of 30 instances show accelerated running time. Specifically, for G11 at $\gamma=0.65$ and G15 at $\gamma=0.85$, the introduction of the pseudo LB technique transforms previously unsolvable instances within the 3-hour limit into solvable ones. For G15 at $\gamma=0.85$, the running time is reduced to 57.88 seconds, corresponding to a speedup of $186\times$. These findings highlight the critical contributions of both preprocessing and pseudo LB techniques in improving efficiency of the proposed method.